# An Empirical Study of Vulnerability Handling Times in CPython

### v1

@@ I Introduction
The questions examined and the paper’s topic in general are easy to motivate. According to benchmarks, Python is the most popular programming language today [7], and because CPython is the most popular interpreter for the language, the vulnerabilities affecting the interpreter affect large user and deployment bases. In addition, as pointed out in the opening Section II, the handling times proxy not only software engineering effort but also security risks. A further motivating point is that the paper’s topic has not been examined previously, despite a large reference literature base on bug and vulnerability handling times, including their fixing times.
@@ III-B2 Independent Variables
    REPORTER: a set of dummy variables identifying reporters of the vulnerabilities. Note that a reporter may or may not be the same person who discovered a given vulnerability. The rationale for the variable builds on existing studies, which have hypothesized and observed that reporters’ and discoverers’ skills and characteristics, including their communication skills, affect bug and vulnerability handling times [3, 10, 17, 26, 34].
@@ V-B Limitations
Also construct validity and robustness of the data might be slightly threatened. As was described in detail in Section III, the presence of two vulnerability tracking systems complicated the data collection process. In addition, many concessions had to be made when operationalizing the variables for the analysis. However, these problems are fairly typical to the research domain. Already the fundamental abstract question of what counts as a vulnerability is not straightforward; a single CVE may reference multiple vulnerabilities and the other way around, at least according to the CPython’s developers.
@@ V-C Further Work
The paper’s main result—the importance of reporters and their characteristics—would deserve a closer look in further research. Why the vulnerability fixing times can be predicted so well by merely identifying the reporters of the CPython vulnerabilities? Although a similar observation has been made also in previous studies, the answers to the question have been only tentative and without much theoretical contributions. If expertise of reporters, including with respect to providing reproducible information, is as important as often seen in the literature [11], it might be also possible to make practical advances by providing better instructions to people about good vulnerability reporting practices. At the moment, the PSF’s policy [21] does not say anything about reproducibility, POCs, and other related things related to vulnerability handling.

### v2

@@ I Introduction
The questions examined and the paper’s topic in general are easy to motivate. According to benchmarks, Python is the most popular programming language today [7], and because CPython is the most popular interpreter for the language, the vulnerabilities affecting the interpreter affect large user and deployment bases. In addition, as pointed out in the opening Section II, the handling times proxy not only software engineering effort but also security risks. A further motivating point is that the paper’s topic has not been examined previously, despite a large reference literature base on bug and vulnerability handling times, including their fixing times. In this regard, the paper contributes to the existing and related work by patching a small but notable knowledge gap.
@@ III-B2 Independent Variables
    REPORTER: a set of dummy variables identifying reporters of the CPython vulnerabilities. Note that a reporter may or may not be the same person who originally discovered a given vulnerability. Reporters were identified by their unique real names or pseudonyms. The rationale for this independent variable builds on many existing studies, which have hypothesized and observed that reporters’ and discoverers’ skills and characteristics, including their communication skills, affect bug and vulnerability handling times [3, 10, 17, 26, 34].
@@ V-B Limitations
Also construct validity and robustness of the data might be slightly threatened. As was described in detail in Section III, the presence of two vulnerability tracking systems complicated the data collection process. In addition, many concessions had to be made when operationalizing the variables for the analysis. However, these problems are fairly typical to the research domain. Already the fundamental abstract question of what counts as a vulnerability is not straightforward; a single CVE may reference multiple vulnerabilities and the other way around, at least according to the CPython’s developers. It should be also acknowledged that manual data collection often involves a degree of subjectivity; in this regard, particularly the POC independent variable may be subjected to criticism.
@@ V-C Further Work
The paper’s main result—the importance of reporters and their characteristics—would deserve a closer look in further research. Why the vulnerability fixing times can be predicted so well by merely identifying the reporters of the CPython vulnerabilities? Although a similar observation has been made also in previous studies, the answers to the question have been only tentative and without much theoretical contributions. If expertise of reporters, including with respect to providing reproducible information, is as important as often seen in the literature [11], it might be also possible to make practical advances by providing better instructions to people about good vulnerability reporting practices. At the moment, the PSF’s policy [21] does not say anything about reproducibility, POCs, and other similar things related to vulnerability handling.